OpenStudy (anonymous):

The greater of two numbers is one less than 8 times the smaller. Their sum is 98. Find the numbers. I always suck at these questions =_=

OpenStudy (anonymous):

It'll be helpful to "let one number be X" Let the smaller number be X. (8X -1) = GREATER NUMBER. Sum of numbers = Smaller + Greater number = X + (8X - 1) = 98 9X - 1 = 98 9X = 99 X = ?

OpenStudy (anonymous):

11